# Towards Robust Pseudo-Label Learning in Semantic Segmentation: An Encoding Perspective

### v1

@@ 1 Introduction
Semantic segmentation has seen significant improvements with recent advances in deep neural networks[long2015fully, chen2017rethinking, cheng2021per, cheng2022masked]. However, a major challenge in semantic segmentation is the requirement of a large volume of fine-grained pixel-level labels, which can be time-consuming and labor-intensive to obtain [cordts2016cityscapes]. Due to the readily available nature of image data, unsupervised domain adaptation (UDA) and semi-supervised learning (SSL) have been introduced in semantic segmentation to handle the label-scarce scenarios. UDA involves learning from synthetic labeled data and transferring knowledge to real unlabeled target domains, while SSL utilizes a tiny portion of annotated data to generalize on unseen data. As a result, UDA and SSL are gaining significant attention as promising approaches to reduce the reliance on extensive annotations in semantic segmentation.
@@
In both UDA and SSL settings, models are trained using annotated and unlabeled data simultaneously. Existing mainstream methods introduce common paradigms, which can be grouped into the self-training pipeline and the consistency regularization framework. Specifically, self-training methods [tranheden2021dacs, jin2022semi] leverage a temporally smoothed exponential moving average (EMA) model as a teacher to generate stable pseudo labels for unlabeled data. On the other hand, the consistency regularization methods [sohn2020fixmatch, araslanov2021self] encourage the model to produce consistent predictions for the same sample across different perturbation views. These paradigms can be summarized as pseudo-label learning, where the network’s predictions are used as supervision for unlabeled data.
@@
Although achieving promising results, the inevitable errors in pseudo labels misled the training process. Typical approaches design filter-out mechanisms [yang2023revisiting, sohn2020fixmatch] and only use high-confidence pseudo-labels for training. However, this paradigm tends to make the model focus on learning from easy samples while neglecting difficult ones, resulting in a sub-optimal performance. Another alternative is to utilize weighting functions [hu2021semi, sun2023daw] that assign weights based on confidence of pseudo-labels. While potentially effective, this approach requires careful design and selection of appropriate hyperparameters, which inevitably compromise its applicability. Based on above discussions, we investigate existing works primarily concentrate on developing specific selection strategies for pseudo-labels but rarely consider the impact of the encoding form assigned for classes. As shown in Fig.1 (a), the pixel features of the class sheep are being confused by the classifier as horse or cow, and an erroneous pseudo-label is typically encoded in a one-hot manner through the argmax operation (see Fig.1 (b)). We speculate that similar classes share common visual attributes, leading to confusing pseudo-labels and further misguiding the training process. How to utilize the shared attributes among confusing classes to design a suitable encoding form for pseudo-label learning is rarely explored.
@@
To explore the encoding form tailored for pseudo-label learning, we explicitly disentangle the classes into fine-grained attributes and consider each class a set of attributes. As shown in Fig.1 (c), even with incorrect predictions in specific attributes, confusing classes still exhibit shared attribute characteristics. For instance, both sheep and cow have horn and hoof and are not biped. Despite potential misclassification, accurate prediction of these shared attributes can still provide valuable guidance for effectively training the network. Based on this observation, we resort to error-correcting output codes (ECOC) [dietterich1994solving] to assign a binary bit string (codeword) as an encoding for each class, decomposing the $N$-class classification problem into $K$ two-class subtasks. The collection of codewords corresponding to each class forms a codebook. This paradigm determines the class by predicting a $K$-bit binary encoding and selecting the nearest neighbor query in the codebook. The encoding form created by suitable ECOC enjoys two properties: class discriminability, ensuring well-separated classes by sufficient Hamming distance between codewords, and attribute diversity, ensured by making each bit-position classifier uncorrelated. ECOC encoding endows the model with the ability to handle partial inaccurate bits and make classification decisions. With a theoretical guarantee (Sec. 4), we show that ECOC can serve as an effective equivalent to one-hot encoding in fully supervised settings, and exhibits greater robustness in pseudo-label learning by achieving a tighter classification error bound under a sufficiently large minimum code distance.
@@ 2.1 Label-scarce Semantic Segmentation
Although deep models have achieved remarkable success in various tasks [krizhevsky2012imagenet, he2016deep, vaswani2017attention, dosovitskiy2020image, chen2024sam, li2023enhancing, wangkai2023maunet, wang2024image], they heavily rely on large amounts of labeled training data and struggle to generalize to data with shifted distributions. This is particularly evident in semantic segmentation, where alternative approaches have been introduced to avoid laborious pixel-wise annotation. Unsupervised domain adaptation (UDA) [ghifary2016deep, ganin2016domain, vu2019advent, toldo2020unsupervised, chen2025alleviate, pan2025exploring, he2025progressive, he2025target, chen2025beyond] aims to transfer knowledge from labeled source domains to unlabeled target domains, enabling models to perform well in the target domain despite distribution differences. Semi-supervised learning (SSL) [chen2021semi, xu2022semi, wang2022semi, zhang2020survey, sun2025towards, sun2025two] leverages a combination of a few manually annotated target samples with a large pool of unlabeled samples to enhance model performance. Weakly supervised learning [wei2016stc, huang2018weakly, ahn2018learning, chan2024scribble] addresses the challenge by utilizing less precise annotation signals, such as image-level labels or bounding boxes. Few-shot learning [dong2018few, wang2019panet, liu2024bidirectional, wu2021learning, li2025dual, luo2024exploring, li2025generalized, li2024localization] approaches tackle scenarios with a small number of annotated samples by leveraging prior knowledge and meta-learning techniques. UDA and SSL, in particular, share a similar objective of using unlabeled (target) data to improve the performance of models trained with labeled (source) data only. In this work, we explore these two settings from a unified perspective of pseudo-label learning.
@@ 2.2 Pseudo-label Learning
Two popular paradigms are often employed when training a model with unlabeled data: self-training-based methods [zou2018unsupervised, tranheden2021dacs, hoyer2022daformer, jin2022semi, li2025balanced] and consistency regularization-based methods [sohn2020fixmatch, chen2021semi, araslanov2021self, yang2023revisiting]. In self-training, the model is trained on unlabeled samples using pseudo labels derived from a teacher network. Consistency regularization aims to ensure prediction stability across different perturbations. Both can be viewed as pseudo-label learning [zheng2021rectifying, lee2013pseudo, qiao2023fuzzy]. More recently, state-of-the-art segmentation methods in UDA and SSL combine both technologies [wang2022freematch, hoyer2023mic, liu2022perturbed, li2025towards_1, sun2025beyond]. Although making significant progress, incorrect pseudo-labels can mislead the model’s training process. Typical approaches incorporate filtering mechanisms [sohn2020fixmatch, yang2023revisiting] to train the model exclusively with highly confident pseudo-labels, while recent research focuses on identifying suitable weight functions to improve training stability [hu2021semi, sun2023daw, hoyer2022daformer]. Some works design new optimization criteria inspired by negative learning [qiao2023fuzzy, wang2022semi] to improve learning from pseudo-labels. In this paper, we focus on the encoding form of the pseudo-label, which is an orthogonal direction to the above approaches.
@@ 4 Theory
In this section, we present our main theoretical results on the performance and robustness of ECOC-based DNNs. We first introduce some key tools and concepts used in our analysis, including the Neural Tangent Kernel (NTK) [jacot2018neural] and its properties in the infinite width limit. Then we state our theorems, which characterize the behavior of ECOC compared to one-hot encoding in the fully supervised setting and the pseudo-label learning setting, respectively. The complete proofs with technical assumptions are deferred to the Appendix A.
@@ 5.1 Experimental Setup
Datasets. We evaluate our approach on two standard benchmarks for synthetic-to-real adaptation of street scenes in the UDA task. The synthetic datasets include GTAv [richter2016playing] (24,966 images) and SYNTHIA [ros2016synthia] (9,400 images). Cityscapes [cordts2016cityscapes], a real-world urban dataset, serves as the target domain, with 2,975 training and 500 validation images. For the SSL setting, we use Cityscapes, PASCAL VOC 2012 [everingham2015pascal], a generic object segmentation benchmark with 1,464 training and 1,449 validation images, along with an augmented set of 10,582 additional training images, and COCO [lin2014microsoft], a challenging benchmark composed of 118k/5k training/validation images with 81 classes.
@@
ProDA [zhang2021prototypical] C 87.8 56.0 79.7 46.3 44.8 45.6 53.5 53.5 88.6 45.2 82.1 70.7 39.2 88.8 45.5 50.4 1.0 48.9 56.4 57.5
CPSL [li2022class] C 92.3 59.5 84.9 45.7 29.7 52.8 61.5 59.5 87.9 41.6 85.0 73.0 35.5 90.4 48.7 73.9 26.3 53.8 53.9 60.8
TransDA [chen2022smoothing] T 94.7 64.2 89.2 48.1 45.8 50.1 60.2 40.8 90.4 50.2 93.7 76.7 47.6 92.5 56.8 60.1 47.6 49.6 55.4 63.9
ADFormer [he2025attention] T 96.7 75.1 88.8 57.5 45.9 45.6 55.4 59.8 90.2 45.6 92.1 70.8 43.0 91.0 78.9 79.3 68.7 52.7 65.0 69.2
CDAC [wang2023cdac] T 97.1 78.7 91.8 59.6 57.1 59.1 66.1 72.2 91.8 53.1 94.5 79.4 51.6 94.6 84.9 87.8 78.7 64.9 67.6 75.3
DACS [tranheden2021dacs] C 89.9 39.7 87.9 39.7 39.5 38.5 46.4 52.8 88.0 44.0 88.8 67.2 35.8 84.5 45.7 50.2 0.2 27.3 34.0 52.1
@@
ProDA [zhang2021prototypical] C 87.8 45.7 84.6 37.1 0.6 44.0 54.6 37.0 88.1 - 84.4 74.2 24.3 88.2 - 51.1 - 40.5 45.6 55.5
CPSL [li2022class] C 87.2 43.9 85.5 33.6 0.3 47.7 57.4 37.2 87.8 - 88.5 79.0 32.0 90.6 - 49.4 - 50.8 59.8 57.9
TransDA [chen2022smoothing] T 90.4 54.8 86.4 31.1 1.7 53.8 61.1 37.1 90.3 - 93.0 71.2 25.3 92.3 - 66.0 - 44.4 49.8 59.3
ADFormer [he2025attention] T 91.8 53.6 87.0 40.5 5.2 46.8 52.1 54.9 88.4 - 92.6 72.5 45.7 86.1 - 61.6 - 50.4 64.4 62.1
CDAC [wang2023cdac] T 93.1 68.5 89.8 51.2 8.9 59.4 65.5 65.3 84.7 - 94.4 81.2 57.0 90.5 - 56.9 - 66.8 66.4 68.7
DACS [tranheden2021dacs] C 80.6 25.1 81.9 21.5 2.9 37.2 22.7 24.0 83.7 - 90.8 67.6 38.3 82.9 - 38.9 - 28.5 47.6 48.3
@@
UDA Setting. We evaluate ECOCSeg on three widely used frameworks, DACS [tranheden2021dacs] with ResNet101 [he2016deep] backbone, DAFormer [hoyer2022daformer], and MIC [hoyer2023mic], with MIT-B5 [xie2021segformer] backbone. Experiments are conducted on one RTX-3090 GPU for DACS and DAFormer, and two for MIC. The network is trained for 40K iterations (batch size 2) using AdamW optimizer with learning rates of $6\times 10^{-5}$ (encoder) and $6\times 10^{-4}$ (decoder), weight decay of 0.01, and linear warm-up for the first 1.5K iterations. Images are rescaled and randomly cropped to $512\times 512$ following DAFormer’s augmentation, and the EMA coefficient for updating the teacher net is 0.999.
@@ 5.3 ECOCSeg for SSL
Table 2: SSL performance on Pascal. The 321321 and $513$ denote the training resolution.
@@
U2PL [wang2022semi] 513 72.0 75.1 76.2 74.4 77.6 78.7
PS-MT [liu2022perturbed] 513 72.8 75.7 76.4 75.5 78.2 78.7
DAW [sun2023daw] 513 76.2 77.6 77.4 78.5 78.9 79.6
RankMatch [mai2024rankmatch] 513 76.6 77.8 78.3 78.9 79.2 80.0
Fixmatch [sohn2020fixmatch] 513 70.6 73.9 75.1 74.3 76.3 76.9
@@ Appendix A Proofs of the Theoretical Results
Lemma A.5 (NTK convergence, [jacot2018neural]).
@@
Lemma A.6 (Hidden Layer Output Bound, [yuerror]).
@@ Appendix B Implementation of Encoding Strategy
Max-min distance encoding. A good error-correcting output code should satisfy row separation and column separation. For the former, the minimum Hamming distance between each pair of codewords $d_{min\_r}$ should be maximized, which can correct at least $\lfloor\frac{d_{min\_r}-1}{2}\rfloor$ single bit errors. As for the latter, each bit classifier should be uncorrelated with the others. This can be ensured by maximizing $d_{min\_c}$ and $N-d_{max\_c}$, which are calculated between columns of the code matrix. Although searching for an optimal encoding matrix is known as an NP-hard problem [crammer2002learnability], we can obtain a sufficiently valuable encoding matrix, denoted as $M_{mmd}$, through a random generation strategy due to the sparsity of the encoding space.
@@
Text-based encoding. According to the above criteria, we can design encoding matrices that exhibit desirable properties and sufficient robustness. However, the resulting codewords are class-agnostic, and the corresponding binary classification problem may be difficult to optimize. Other than manually designing encodings based on class attributes, we adopt a concise and automated method based on text embedding to generate codewords for the classes. Specifically, $N$ class names are mapped to the feature space $f_{text}\in\mathbb{R}^{N\times C}$ through word2vec [mikolov2013distributed]. Then, we compress extracted continuous features and quantize them into binary encodings of length $K$ to obtain the code matrix, denoted as $M_{text}$. This encoding strategy considers the relationships and structural information among classes, facilitating more efficient encoding learning.
To consider the relationships and structural information among classes, we resort to word2vec [mikolov2013distributed] to extract class-related features. Then, we select the most discriminant feature components based on variance magnitude to compress the feature dimensionality to the length of the codeword. Furthermore, we quantize the features into 0-1 encoding using the mean of the feature components as a threshold. During this process, we also need to ensure the validity of the codebook, as discussed above.

### v2

@@ 1 Introduction
Semantic segmentation has seen significant improvements with recent advances in deep neural networks[55, 9, 17, 16]. However, a major challenge in semantic segmentation is the requirement of a large volume of fine-grained pixel-level labels, which can be time-consuming and labor-intensive to obtain [19]. Due to the readily available nature of image data, unsupervised domain adaptation (UDA) and semi-supervised learning (SSL) have been introduced in semantic segmentation to handle the label-scarce scenarios. UDA involves learning from synthetic labeled data and transferring knowledge to real unlabeled target domains, while SSL utilizes a tiny portion of annotated data to generalize on unseen data. As a result, UDA and SSL are gaining significant attention as promising approaches to reduce the reliance on extensive annotations in semantic segmentation.
@@
In both UDA and SSL settings, models are trained using annotated and unlabeled data simultaneously. Existing mainstream methods introduce common paradigms, which can be grouped into the self-training pipeline and the consistency regularization framework. Specifically, self-training methods [79, 41] leverage a temporally smoothed exponential moving average (EMA) model as a teacher to generate stable pseudo labels for unlabeled data. On the other hand, the consistency regularization methods [72, 3] encourage the model to produce consistent predictions for the same sample across different perturbation views. These paradigms can be summarized as pseudo-label learning, where the network’s predictions are used as supervision for unlabeled data.
@@
Although achieving promising results, the inevitable errors in pseudo labels misled the training process. Typical approaches design filter-out mechanisms [93, 72] and only use high-confidence pseudo-labels for training. However, this paradigm tends to make the model focus on learning from easy samples while neglecting difficult ones, resulting in a sub-optimal performance. Another alternative is to utilize weighting functions [36, 77] that assign weights based on confidence of pseudo-labels. While potentially effective, this approach requires careful design and selection of appropriate hyperparameters, which inevitably compromise its applicability. Based on above discussions, we investigate existing works primarily concentrate on developing specific selection strategies for pseudo-labels but rarely consider the impact of the encoding form assigned for classes. As shown in Fig.1 (a), the pixel features of the class sheep are being confused by the classifier as horse or cow, and an erroneous pseudo-label is typically encoded in a one-hot manner through the argmax operation (see Fig.1 (b)). We speculate that similar classes share common visual attributes, leading to confusing pseudo-labels and further misguiding the training process. How to utilize the shared attributes among confusing classes to design a suitable encoding form for pseudo-label learning is rarely explored.
@@
To explore the encoding form tailored for pseudo-label learning, we explicitly disentangle the classes into fine-grained attributes and consider each class a set of attributes. As shown in Fig.1 (c), even with incorrect predictions in specific attributes, confusing classes still exhibit shared attribute characteristics. For instance, both sheep and cow have horn and hoof and are not biped. Despite potential misclassification, accurate prediction of these shared attributes can still provide valuable guidance for effectively training the network. Based on this observation, we resort to error-correcting output codes (ECOC) [21] to assign a binary bit string (codeword) as an encoding for each class, decomposing the $N$-class classification problem into $K$ two-class subtasks. The collection of codewords corresponding to each class forms a codebook. This paradigm determines the class by predicting a $K$-bit binary encoding and selecting the nearest neighbor query in the codebook. The encoding form created by suitable ECOC enjoys two properties: class discriminability, ensuring well-separated classes by sufficient Hamming distance between codewords, and attribute diversity, ensured by making each bit-position classifier uncorrelated. ECOC encoding endows the model with the ability to handle partial inaccurate bits and make classification decisions. With a theoretical guarantee (Sec. 4), we show that ECOC can serve as an effective equivalent to one-hot encoding in fully supervised settings, and exhibits greater robustness in pseudo-label learning by achieving a tighter classification error bound under a sufficiently large minimum code distance.
@@ 2.1 Label-scarce Semantic Segmentation
Although deep models have achieved remarkable success in various tasks [42, 29, 80, 23, 13, 48, 88, 86], they heavily rely on large amounts of labeled training data and struggle to generalize to data with shifted distributions. This is particularly evident in semantic segmentation, where alternative approaches have been introduced to avoid laborious pixel-wise annotation. Unsupervised domain adaptation (UDA) [27, 26, 81, 78, 15, 64, 32, 31, 14] aims to transfer knowledge from labeled source domains to unlabeled target domains, enabling models to perform well in the target domain despite distribution differences. Semi-supervised learning (SSL) [12, 92, 87, 99, 76, 75] leverages a combination of a few manually annotated target samples with a large pool of unlabeled samples to enhance model performance. Weakly supervised learning [89, 37, 1, 6] addresses the challenge by utilizing less precise annotation signals, such as image-level labels or bounding boxes. Few-shot learning [22, 84, 53, 90, 50, 56, 51, 49] approaches tackle scenarios with a small number of annotated samples by leveraging prior knowledge and meta-learning techniques. UDA and SSL, in particular, share a similar objective of using unlabeled (target) data to improve the performance of models trained with labeled (source) data only. In this work, we explore these two settings from a unified perspective of pseudo-label learning.
@@ 2.2 Pseudo-label Learning
Two popular paradigms are often employed when training a model with unlabeled data: self-training-based methods [105, 79, 34, 41, 46] and consistency regularization-based methods [72, 12, 3, 93]. In self-training, the model is trained on unlabeled samples using pseudo labels derived from a teacher network. Consistency regularization aims to ensure prediction stability across different perturbations. Both can be viewed as pseudo-label learning [103, 44, 66]. More recently, state-of-the-art segmentation methods in UDA and SSL combine both technologies [85, 35, 54, 47, 74]. Although making significant progress, incorrect pseudo-labels can mislead the model’s training process. Typical approaches incorporate filtering mechanisms [72, 93] to train the model exclusively with highly confident pseudo-labels, while recent research focuses on identifying suitable weight functions to improve training stability [36, 77, 34]. Some works design new optimization criteria inspired by negative learning [66, 87] to improve learning from pseudo-labels. In this paper, we focus on the encoding form of the pseudo-label, which is an orthogonal direction to the above approaches.
@@ 4 Theory
In this section, we present our main theoretical results on the performance and robustness of ECOC-based DNNs. We first introduce some key tools and concepts used in our analysis, including the Neural Tangent Kernel (NTK) [38] and its properties in the infinite width limit. Then we state our theorems, which characterize the behavior of ECOC compared to one-hot encoding in the fully supervised setting and the pseudo-label learning setting, respectively. The complete proofs with technical assumptions are deferred to the Appendix A.
@@ 5.1 Experimental Setup
Datasets. We evaluate our approach on two standard benchmarks for synthetic-to-real adaptation of street scenes in the UDA task. The synthetic datasets include GTAv [67] (24,966 images) and SYNTHIA [69] (9,400 images). Cityscapes [19], a real-world urban dataset, serves as the target domain, with 2,975 training and 500 validation images. For the SSL setting, we use Cityscapes, PASCAL VOC 2012 [25], a generic object segmentation benchmark with 1,464 training and 1,449 validation images, along with an augmented set of 10,582 additional training images, and COCO [52], a challenging benchmark composed of 118k/5k training/validation images with 81 classes.
@@
ProDA [100] C 87.8 56.0 79.7 46.3 44.8 45.6 53.5 53.5 88.6 45.2 82.1 70.7 39.2 88.8 45.5 50.4 1.0 48.9 56.4 57.5
CPSL [45] C 92.3 59.5 84.9 45.7 29.7 52.8 61.5 59.5 87.9 41.6 85.0 73.0 35.5 90.4 48.7 73.9 26.3 53.8 53.9 60.8
TransDA [11] T 94.7 64.2 89.2 48.1 45.8 50.1 60.2 40.8 90.4 50.2 93.7 76.7 47.6 92.5 56.8 60.1 47.6 49.6 55.4 63.9
ADFormer [30] T 96.7 75.1 88.8 57.5 45.9 45.6 55.4 59.8 90.2 45.6 92.1 70.8 43.0 91.0 78.9 79.3 68.7 52.7 65.0 69.2
CDAC [83] T 97.1 78.7 91.8 59.6 57.1 59.1 66.1 72.2 91.8 53.1 94.5 79.4 51.6 94.6 84.9 87.8 78.7 64.9 67.6 75.3
DACS [79] C 89.9 39.7 87.9 39.7 39.5 38.5 46.4 52.8 88.0 44.0 88.8 67.2 35.8 84.5 45.7 50.2 0.2 27.3 34.0 52.1
@@
ProDA [100] C 87.8 45.7 84.6 37.1 0.6 44.0 54.6 37.0 88.1 - 84.4 74.2 24.3 88.2 - 51.1 - 40.5 45.6 55.5
CPSL [45] C 87.2 43.9 85.5 33.6 0.3 47.7 57.4 37.2 87.8 - 88.5 79.0 32.0 90.6 - 49.4 - 50.8 59.8 57.9
TransDA [11] T 90.4 54.8 86.4 31.1 1.7 53.8 61.1 37.1 90.3 - 93.0 71.2 25.3 92.3 - 66.0 - 44.4 49.8 59.3
ADFormer [30] T 91.8 53.6 87.0 40.5 5.2 46.8 52.1 54.9 88.4 - 92.6 72.5 45.7 86.1 - 61.6 - 50.4 64.4 62.1
CDAC [83] T 93.1 68.5 89.8 51.2 8.9 59.4 65.5 65.3 84.7 - 94.4 81.2 57.0 90.5 - 56.9 - 66.8 66.4 68.7
DACS [79] C 80.6 25.1 81.9 21.5 2.9 37.2 22.7 24.0 83.7 - 90.8 67.6 38.3 82.9 - 38.9 - 28.5 47.6 48.3
@@
UDA Setting. We evaluate ECOCSeg on three widely used frameworks, DACS [79] with ResNet101 [29] backbone, DAFormer [34], and MIC [35], with MIT-B5 [91] backbone. Experiments are conducted on one RTX-3090 GPU for DACS and DAFormer, and two for MIC. The network is trained for 40K iterations (batch size 2) using AdamW optimizer with learning rates of $6\times 10^{-5}$ (encoder) and $6\times 10^{-4}$ (decoder), weight decay of 0.01, and linear warm-up for the first 1.5K iterations. Images are rescaled and randomly cropped to $512\times 512$ following DAFormer’s augmentation, and the EMA coefficient for updating the teacher net is 0.999.
@@ 5.3 ECOCSeg for SSL
Table 2: SSL performance on Pascal. The $321$ and $513$ denote the training resolution.
@@
U2PL [87] 513 72.0 75.1 76.2 74.4 77.6 78.7
PS-MT [54] 513 72.8 75.7 76.4 75.5 78.2 78.7
DAW [77] 513 76.2 77.6 77.4 78.5 78.9 79.6
RankMatch [58] 513 76.6 77.8 78.3 78.9 79.2 80.0
Fixmatch [72] 513 70.6 73.9 75.1 74.3 76.3 76.9
@@ Appendix A Proofs of the Theoretical Results
Lemma A.5 (NTK convergence, [38]).
@@
Lemma A.6 (Hidden Layer Output Bound, [96]).
@@ Appendix B Implementation of Encoding Strategy
Max-min distance encoding. A good error-correcting output code should satisfy row separation and column separation. For the former, the minimum Hamming distance between each pair of codewords $d_{min\_r}$ should be maximized, which can correct at least $\lfloor\frac{d_{min\_r}-1}{2}\rfloor$ single bit errors. As for the latter, each bit classifier should be uncorrelated with the others. This can be ensured by maximizing $d_{min\_c}$ and $N-d_{max\_c}$, which are calculated between columns of the code matrix. Although searching for an optimal encoding matrix is known as an NP-hard problem [20], we can obtain a sufficiently valuable encoding matrix, denoted as $M_{mmd}$, through a random generation strategy due to the sparsity of the encoding space.
@@
Text-based encoding. According to the above criteria, we can design encoding matrices that exhibit desirable properties and sufficient robustness. However, the resulting codewords are class-agnostic, and the corresponding binary classification problem may be difficult to optimize. Other than manually designing encodings based on class attributes, we adopt a concise and automated method based on text embedding to generate codewords for the classes. Specifically, $N$ class names are mapped to the feature space $f_{text}\in\mathbb{R}^{N\times C}$ through word2vec [60]. Then, we compress extracted continuous features and quantize them into binary encodings of length $K$ to obtain the code matrix, denoted as $M_{text}$. This encoding strategy considers the relationships and structural information among classes, facilitating more efficient encoding learning.
To consider the relationships and structural information among classes, we resort to word2vec [60] to extract class-related features. Then, we select the most discriminant feature components based on variance magnitude to compress the feature dimensionality to the length of the codeword. Furthermore, we quantize the features into 0-1 encoding using the mean of the feature components as a threshold. During this process, we also need to ensure the validity of the codebook, as discussed above.